OpenStudy (anonymous):

if f(x)=3x+9, what is f^-1(x)

OpenStudy (anonymous):

Do you know how to find the inverse of an equation?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

y=3x+9 swap places for x and y, then solve for y x=3y+9 y=(x-9)/3

OpenStudy (anonymous):

f^-1(x)=(1/3)x-3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yep